OpenStudy (anonymous):

Evaluate the following limits as x approaches 0. i. x/ln (1+x) ii. xx

OpenStudy (anonymous):

for i., use l'Hopital's Rule... for ii., why don't you just substitute x=0 because x*x=x^2 which is continuous over all x.

OpenStudy (anonymous):

thanks i had 1/(1+x)

OpenStudy (anonymous):

for i? isn't it 1/(1/(1+x)) = 1+x ?

OpenStudy (turingtest):

is the second one\[x\cdot x\]or\[x^x\]?

OpenStudy (anonymous):

\[x ^{x}\]